# Image Matching Filtering and Refinement by Planes and Beyond

### v3

@@ 4.2.9 Visual qualitative analysis
+MAGSAC↑ |
@@ 4.2.10 Running times
In general, the base matching pipelines run within 0.5 s with exception of SIFT and Key.Net, which run within 1 s since their pipelines are not completely based on PyTorch and do not exploit the inherent parallelization offered. The overhead added by the filtering modules is almost irrelevant for all the compared filters except CC and the proposed ones. Specifically, CC lacks a parallel implementation and can roughly process 1000 matches in 3 s, NNC is able to refine about 3000 patch pairs within 1 s, while MOP and MOP+MiHo computations depend not only on the size of the input but also on the outlier rate of the input images. In this latter case, running times can be mainly controlled by the number of the maximum iterations for RANSAC , defined in Sec. 3.1.3 and set to 500, 1000, 1500 and 2000 for the analysis. It is not trivial to derive the running time trend for MOP and MOP+MiHo. In general, MOP goes roughly from 1 s to 2 s for SIFT and Key.Net moving from 500 to 2000, and from 2.5 s to 7s with other pipelines. For MOP+MiHo, setting from 500 to 2000 changes the running times from 1.5 s to 4 s for all pipelines except LoFTR and DeDoDe v2, for which the running time moves from 3 s to 6 s. Compared with other filtering and refinement modules MOP and MOP+MiHo running time is relatively high but still acceptable for off-line, medium-scale image matching, as well as for the NCC refinement module. The proposed modules have been implemented in PyTorch exploiting its base parallel support and both GPU workload and memory usage were around 25% when processing data. Likely, hardware-related and coding optimization could speed up these modules. Notice also that deep-architecture is parallel by design while MOP-based computation inherits RANSAC serialization. Nevertheless, better optimization to adaptively determine the stopping criteria and to avoid redundant computation dynamically or in a coarse-to-fine manner are planned as future work.
Overall, according to the running time data, it turns out that the default value =2000 used for the main comparative evaluation in Tables 1-7 is optimal in terms of matching quality only for SIFT and Key.Net, while it slightly decreases the matching results for the other pipelines with respect to =500, which also improves the running times. In detail, moving from =500 to =2000 causes an increment of around 2% of the recall and of about 1% for the AUC pose estimation in the case of SIFT and Key.Net, which contain as initial configuration a higher number of outliers than other pipelines, for which both recall and pose AUC remain instead stable or even worsen. This can be motivated by observing that the next homography discovered in MOP-based approaches is more noisy or even not correct in the long run than the previous ones, and high values force to retrieve a plane and delay the iteration stops. According to these considerations, it is reasonable that forcing the homography retrieval by increasing starts to degrade the found solutions as the initial outlier ratio decreases.
Notice also that MiHo can take part in the MOP early stopping since it introduces further constraints in the homography to be searched, justifying higher running times for MOP than those of MOP+MiHo in some configurations, although MiHo should ideally need more computational resources as two tied homographies are estimated at each step instead of only one. In particular, MOP+MiHo running time equals MOP one for LoFTR and DeDoDe v2, it is greater for SuperPoint, ALIKED, and DISK, and it is lower for SIFT and Key.Net. Excluding SIFT and Key.Net, MOP+MiHo find more homographies than MOP, suggesting the reliability of the organization of input matches within the scene as almost-similarity transformations.
@@ 5 Conclusions and future work
MOP+MiHo+NCC is able to effectively discard outliers and to improve the match localization as well as the pose accuracy with Key.Net, SuperPoint, and LoFTR. MOP+MiHo without NCC filtering is able to remove outliers within SIFT providing a better RANSAC estimation as well. In the case of ALIKED, DISK, and DeDoDe v2, where the base pipeline only contains accurate matches, the proposed pipeline does not alter the final output. Nevertheless, in terms of absolute accuracy, LoFTR and SuperPoint reached the overall best results among all the evaluated pipelines when boosted by the proposed approach in three of the evaluated datasets, while in the remaining IMC-PT dataset, the best is reached by DISK with FC-GNN, which is the current deep competitor and alternative to MOP+MiHo+NCC.
Overall, the above behavior suggests that current end-to-end deep base matching architecture are in general too strict in removing non-accurate matches which could be refined. In this sense, the proposed analysis and the designed approaches provide further insight that can be used for developing better deep image matching networks. FC-GNN provides better pose estimation than the proposed method for unknown camera intrinsics, with a minimal gap with respect to the pose estimation with known camera intrinsics in MegaDepth. Future research will involve the inclusion of camera intrinsic statistics and heuristics within the RANSAC framework.
MOP has proved to be effective in finding spatial planar clusters. As future work these clusters could be organized and exploited more in-depth to further filter matches as well as to expand them within the framework described in [50]. Moreover, planar clusters could be used also to characterize the scene structure at higher levels in terms of object segmentation and the relative positions and occlusions of the found elements.
@@
NCC inherits the strengths and weaknesses of the base template matching approach it is based on. NCC is effective on corner-like patches but on strong edge-like structures or flat homogeneous regions, more typical of blobs, NCC can mislead the alignment refinement. More elaborated and robust approaches based on the gradient orientation histogram correlation, on the Iterative Closest Point (ICP) [110] or variational methods will be taken in consideration for future developments. Match densification through local patch homographies will be investigated too.

### v4

@@ 4.2.7 Recent dense pipelines
4.2.8 Visual qualitative analysis
Figure 6 shows visual results of the patch alignment by NCC for two example image pairs, superimposed as red-green anaglyphs. Patches have been randomly chosen for visualization, and the full patch sets for each evaluated pipeline and dataset combination can be generated by the provided supporting tool1. Inside each multirow in the figure, the top row shows the initial patch alignment specified by the homography pair defined in Sec. 3.3, while the second row shows the results after applying NCC without MOP-based homography solutions. Likewise, the third row shows the initial alignment of the MOP+MiHo extended homography pair and the bottom row the final results with NCC. In the left clusters , with or without NCC, has the highest correlation value. The blue bar on the left of each patch superimposition indicates the corresponding epipolar error of the patch center with respect to the pseudo GT. The reader is invited to zoom in and inspect the above figure, which provides examples of the possible configurations and alignment failure cases.
For true planar regions, such as the patches including the St. Peter’s Basilica epigraph, it is evident that MOP+MiHo extended patches are better than the base ones in terms of alignment and distortions errors with respect to the original images. Nevertheless, MOP patch alignment can go wrong in some circumstances as shown in the leftmost column of the top multirow. This happens for instance when the homography assignment described in Sec. 3.1.4 picks a homography with a small reprojection error but also a poor wrong support set. Furthermore, NCC similarity maximization is not always the best choice as highlighted by the patches containing the inscription in the left clusters, where the registration is visibly better with MOP although base patches achieve the highest NCC correlation value.
In the case of patches representing almost flat regions or straight contour edges according to the Harris corner definition [harris], NCC can be misled and can degrade the alignment increasing the epipolar error, as in the rightmost patch overlay in the third multirow. Notice also that although visually improving the overall patch alignment, NCC can increase the epipolar error of the effective keypoint center, as for the ‘BV’ inscription in the third multirow. Anyways, when NCC keypoint accuracy decreases assuming that the pseudo-GT sub-pixel estimation can be always trusted, the epipolar error increment is generally limited as reported in the figure and by the results of Tables 1-5.
@@ 4.2.8 Visual qualitative analysis
+MAGSAC |
Figures 7-8 present some example image pairs, respectively in the case of non-planar and planar scenes, for base matching pipelines where NCC refinement is effective according to the quantitative evaluation. The base matches after being filtered by MOP+MiHo+NCC and FC-GNN, with or without RANSAC, are reported, where the match color indicates the corresponding maximum epipolar error within the pair images.
Comparing the first row for each scene, i.e. without RANSAC-based filtering, both MOP+MiHo and FC-GNN remove outliers, highlighted in blue, but MOP+MiHo is more relaxed in the task, acting as a sort of coarse Local Optimized RANSAC (LO-RANSAC) [lo-ransac] step. As reported in Tables 1-5, this is beneficial for the effective RANSAC especially when the outlier contamination of the input matches is high, such as for SIFT and Key.Net, and does not negatively affect the final results in the other cases. Concerning the refinement, both NCC and FC-GNN are able to improve the keypoint accuracy, as indicated by the color shift of the matches, but the epipolar error variation of their refined matches differs. For NCC the same observations and considerations discussed for Figure 6 hold.
+MOP+MiHo+NCC |
+FC-GNN |
| MASt3R | ![]() |
+MAGSAC |
| MASt3R | ![]() |
+MAGSAC |
Finally, Fig. 9 presents some example image pairs in the case of non-planar and planar scenes when MASt3R is the base pipeline, with the aim to complete the observations presented in Sec. 4.2.7. The MASt3R base matches are not very accurate in terms of maximum epipolar error as indicated by the red and purple colored matches in the left column, but these are quite consistent with the true epipolar geometry so these are almost all preserved by MAGSAC. In the center column, MOP does not remove ambiguous matches in the sky which are globally consistent with a virtual plane, while NCC effectively increases the accuracy for most of the matches. Nevertheless, in the case where the are no real matches such as for the fountain and crowd in the bottom part of the non-planar scene image, NCC provides non-matches with higher epipolar error which are no more globally consistent with the real true scene pose. Moreover, NCC fails in the bottom right part of the arcade due to edge-like patches not feasible for template matching. Thus the pose estimation becomes less accurate. This does not happen in the planar scenes due to the different nature of the problem and of the constraints so that NCC provides a remarkable improvement. In the last column, one can see that FC-GNN in also able to improve the precision of the matches but it obtains more yet less accurate matches than MOP+MiHo+NCC. Differently from the proposed strategy, both MASt3R and FC-GNN are trained deep models relying on epipolar pose priors.
4.2.9 Ablation study
Table 6 reports the results obtained using several variations of the MOP-based approaches, including the reflection check through MOP⋆ (see Sec. 3.2.5), the replacement of MiHo with VSAC (see Sec. 3.2.3), and taking into account in template matching the patch gradient orientation through NCC⋆ (see Sec. 3.3). The results are reported for the base pipelines where MOP-based approaches are effective and, besides the base pipeline, include FC-GNN which is the best overall deep alternative. Extended results are available as additional material6.
Concerning MOP and MOP⋆ there are basically no differences. As there are no real planar reflections in the datasets, this confirms the robustness of the process, and the small fluctuations are likely due to the randomness of the MOP inner RANSAC process.
NCC⋆ usually provides an improvement upon NCC in terms of accuracy around 0.5-1%, especially for the fundamental matrix estimation. NCC⋆ can be slightly worse in some cases, but even in the most noticeable cases, i.e. the essential matrix estimation for MegaDepth and the planar homography case for both SuperPoint and Key.Net, the degradation is quite negligible.
VSAC on average is slightly worse than MiHo, the differences between the two approaches are minimal. Only in the planar homography estimation VSAC is slightly better than MiHo, probably due to the fact that VSAC tends to adhere to the strict planar homography model more than MiHo.
Overall, MOP⋆+MiHo+NCC⋆ seem to provide the best combination, even if compared with MOP+MiHo+NCC these are not impressive. With respect to the base pipeline the MOP-based approaches provide evident improvements for the non-planar datasets, while on the planar dataset the differences are minimal with the exception of LoFTR which benefits of the filtering and refinements of the matches.
Compared against the deep FC-GNN, MOP-based combinations with NCC offer a valid alternative, being the results comparable or better. Notice that FC-GNN is effectively preferable only for the fundamental estimation on MegaDepth were it was trained on, and in the planar dataset is noticeably worse than the base pipeline.
Note also that even for the base pipelines presented in Sec. 4.2.6 and Sec. 4.2.7, the degradation of MOP-based filtering is almost null compared with other filtering approaches. The only exception is for the non-planar scenes with MASt3R when NCC is added to MOP-base filters for the reasons explained in Sec. 4.2.7. This implies that the approach presented is robust and stable.
@@ 4.2.10 Running times
The base matching pipelines run within 0.5-1 s in the testing environment employed. The overhead added by the filtering modules is almost irrelevant for all the compared filters except CC and the proposed ones. NCC is able to refine about 3000 patch pairs within 1 s, while MOP and MOP+MiHo computations depend not only on the size of the input but also on the outlier rate of the input images. In this latter case, running times can be mainly controlled by the number of the maximum iterations for RANSAC , defined in Sec. 3.1.3 and set to 500, 1000, 1500 and 2000 for the analysis. It is not trivial to derive the running time trend for MOP and MOP+MiHo. In general, MOP goes roughly from 1 s to 2 s for SIFT and Key.Net moving from 500 to 2000, and from 2.5 s to 7s with other pipelines. For MOP+MiHo, setting from 500 to 2000 changes the running times from 1.5 s to 4 s for all pipelines except LoFTR and DeDoDe v2, for which the running time moves from 3 s to 6 s. More details can be found in Appendix E.5.
Compared with other filtering and refinement modules, MOP and MOP+MiHo running times are relatively high but still acceptable for off-line, medium-scale image matching tasks. The same holds for the running times of the NCC refinement module. The proposed modules have been implemented in PyTorch exploiting its base parallel support and both GPU workload and memory usage were around 25% when processing data. Likely, hardware-related and coding optimization could speed up these modules. Notice also that deep-architecture are parallel by design while MOP-based computation inherits RANSAC serialization. Nevertheless, better optimizations to adaptively determine the stopping criteria and to avoid redundant computation dynamically or in a coarse-to-fine manner are planned as future work.
@@ 5 Conclusions and future work
MOP+MiHo+NCC and its variants are able to effectively discard outliers and to improve the match localization as well as the pose accuracy with SuperPoint, LoFTR, Key.Net and SIFT. When the base pipeline only contains accurate matches such as for ALIKED, DISK, and DeDoDe v2 and RoMa, the proposed pipeline does not alter the final output being stable and robust. In terms of absolute accuracy, LoFTR and SuperPoint reached overall best results among all the evaluated sparse pipelines when boosted by the proposed approach for MegaDepth and ScanNet. For the remaining non-planar IMC-PT dataset, the best is reached by DISK with FC-GNN, which is the current deep competitor and alternative to MOP+MiHo+NCC.
Overall, the above behavior suggests that current sparse end-to-end deep base matching architecture are in general too strict in removing non-accurate matches which could be refined. In this sense, the proposed analysis and the designed approaches provide further insight that can be used for developing better deep image matching networks.
MOP has proved to be effective in finding spatial planar clusters. As future work these clusters could be organized and exploited more in-depth to further filter matches as well as to expand them within the framework described in [slime], also taking into account the epipolar geometry of the scene. Moreover, planar clusters could be used also to characterize the scene structure at higher levels in terms of object segmentation and the relative positions and occlusions of the found elements.
@@
NCC inherits the strengths and weaknesses of the base template matching approach it is based on. NCC is effective on corner-like patches but on strong edge-like structures or flat homogeneous regions, more typical of blobs, NCC can mislead the alignment refinement. Nevertheless, NCC⋆ shows that base template matching can be improved exploiting further options. More elaborated and robust approaches based on the gradient orientation histogram correlation, on the Iterative Closest Point (ICP) [icp] or variational methods will be taken in consideration for future developments. Match densification through local patch homographies will be investigated too.
@@
Declarations
Data Availability Statement All the resources used within this paper, including code and data, are freely available at the project research page https://github.com/fb82/MiHo.
Author contribution F. Bellavia: Conceptualization, Methodology, Software, Validation, Investigation, Resources, Data Curation, Writing - Original Draft, Writing - Review & Editing, Visualization, Supervision; Z.Zhao: Software, Investigation, Resources, Writing - Review & Editing, Visualization; F. Remondino: Investigation, Resources; L. Morelli: Investigation, Resources.
Appendix A MiHo rotation handling
A simple strategy was experimentally verified to work well in practice for dealing with rotations when employing MiHo, as discussed in Sec. 3.2.5, under the observation that images acquired by cameras tend to have a possible relative rotation of a multiple of .
where
and is the Iverson bracket. The global orientation estimation can be computed efficiently on the initial input matches before running MOP, adjusting keypoints accordingly. Final homographies can be adjusted in turn by removing the rotation given by . Figure 10 shows the MOP+MiHo result obtained on a set of matches without or with orientation pre-processing, highlighting the better matches obtained in the latter case. This approach can easily be extended to handle image reflections.
Appendix B Check for reflections
As discussed in Sec. 3.2.5, a further check can be used directly inside the RANSAC loop in Sec. 3.1.3 to further constraint the planar search when no image reflections are allowed as for the datasets used in the evaluation. This further variant of MOP is denoted by MOP⋆.
Given the minimum hypothesis sampled model of Eq. 9, it is possible to define four triangles in the image , according to the triplets
where . There are no reflections if the corresponding triangles and considering the ordering provided by the triplet maintains the same clockwise or counterclockwise orientation. Hence it must holds that
where
Notice that although similar in terms of approach, the reflection check is different from the twisted quadrilateral condition often used as homography constraint [superransac], implicitly taken into account in MOP RANSAC implementation by the quasi-affinity check described in Sec. 3.1.3.
Both MOP and MOP⋆ uses the same fixed random seed for comparative purposes.
Appendix C NCC refinement
As presented in Sec. 3.3, the keypoint refinement is achieved by obtaining the pixel offset which maximizes the NCC similarity of the corresponding patches among all the warping homography pairs contained in the set , defined in Eq. 3.3. Finally, parabolic interpolation is used to get a sub-pixel refinement.
C.1 NCC computation
such that the squared window set of offsets is
and has cardinality equal to
The squared patch of centered in with radius is the pixel set
while the mean intensity value of the patch is
and the variance is
which is ideally robust to affine illumination changes [gonzales_wood]. Lastly, the similarity between two patches and is
so that, for the match and a set of associated homography pairs derived in Eq. 3.3, the refined keypoint offsets and the best aligning homography pair are given by
is obtained by reprojecting back to the original images, i.e.
C.2 Sub-pixel precision
The refinement offset can be further enhanced by parabolic interpolation adding sub-pixel precision [trucco_verri]. Notice that other forms of interpolation have been investigated but they did not provide any sub-pixel accuracy improvements according to previous experiments [miho_base]. The NCC response map at in the best warped space with the origin coinciding with the peak value can be written as
from computed by Eq. C.1. The sub-pixel refinement offsets in the horizontal direction are computed as the vertex of a parabola fitted on the horizontal neighborhood centered in the peak of the maximized NCC response map, i.e.
and from Eq. C.1 the final refined match is
with
Appendix D Evaluation details
D.1 Base pipelines
As introduced in Sec. 4.1.1, there are seven base pipelines employed for the main comparative analysis carried out in the paper. With the exception of LoFTR, all the evaluated methods are sparse image matching approaches.
SIFT+NNR [sift] is included as the standard and reference handcrafted pipeline. The OpenCV [opencv] implementation was used for SIFT, exploiting RootSIFT [rootsift] for descriptor normalization, while NNR implementation is provided by Kornia [kornia]. To stress the match filtering robustness of the successive steps of the pipeline, which is the goal of this evaluation, the NNR ratio threshold was set rather high, i.e. to 0.95. Common adopted values range in depending on the scene complexity, with higher values yielding to less discriminative matches and then possible outliers. Moreover, upright patches are employed by zeroing the dominant orientation for a fair comparison with other recent deep approaches.
Key.Net++NNR [imw2020], a modular pipeline that achieves good results in common benchmarks, is also taken into account. Excluding the NNR stage, it can be considered a modular deep pipeline. The Kornia implementation is used for the evaluation. As for SIFT, the NNR threshold is set very high, i.e. to 0.99, while more common values range in . The deep orientation estimation module OriNet [affnet] generally siding AffNet is not included to provide upright matches.
Other base pipelines considered are SOTA full end-to-end matching networks. In details, these are SuperPoint+LightGlue, DISK+LightGlue and ALIKED+LightGlue implemented by [lightglue]. The input images are rescaled so the smaller size is 1024, following LightGlue default. For ALIKED the aliked-n16rot model weights are employed which according to the authors [aliked] can handle slight image rotation better and are more stable on viewpoint changes.
The last base pipelines added in the evaluation are DeDoDe v2 [dedodev2], which provides an alternative end-to-end deep architecture different from the above, and LoFTR, a semi-dense detector-free end-to-end network. These latter pipelines achieve SOTA results in popular benchmarks. The authors’ implementation is employed for DeDoDe v2, setting to matching threshold to 0.01, while LoFTR implementation available through Kornia is used in the latter case.
Further experiments on a more selective evaluation configurations have been done using the recent RoMa [mast3r] and MAST3R [roma] dense architectures, leading to a total of nine evaluated pipelines. Both RoMa and MASt3R achieve impressive results, especially in complex scenarios. The implementations of their respective authors have been used. For RoMa the coarse and upsample resolutions were set respectively to 560 px and 864 px, while for MASt3R the input images were resized to 512 px. In both case the maximum number of matches was limited to 2048, considering computational resource constraints, and taking into account the low number of returned outliers.
D.2 Datasets
As discussed in Sec. 4.1.3, evaluations have been carried out using four different datasets. The dataset statistics employed to extract the intrinsic camera matrix parameter statistics described in Sec. 4.1.2 are shown in Fig. 12.
MegaDepth and ScanNet, respectively outdoor and indoor datasets, are de-facto standard in nowadays image matching benchmarking; sample image pairs for each dataset are shown in Fig. 1. For these datasets, the same 1500 image pairs for each dataset and GT data indicated in the protocol employed in [loftr] are used. MegaDepth test image pairs belong to only 2 different scenes and are resized proportionally so that the maximum size is equal to 1200 px, while ScanNet image pairs belong to 100 different scenes and are resized to px. Although according to previous evaluation [lightglue] LightGlue provides better results when the images are rescaled so that the maximum dimension is 1600 px, in this evaluation the original 1200 px resize was maintained since more outliers and minor keypoint precision are achieved in the latter case, providing a configuration in which the filtering and refinement of the matches can be better revealed. For deep methods, the best weights that fit the kind of scene, i.e. indoor or outdoor, are used when available.
The IMC-PT dataset is a curated collection of sixteen scenes derived by the SfM models of the Yahoo Flickr Creative Commons 100 Million (YFCC100M) dataset [yfcc100m]. For each scene 800 image pairs have been randomly chosen, resulting in a total of 12800 image pairs. These scenes also provide 3D model scale as GT data so that metric evaluations are possible. Note that MegaDepth data are roughly a subset of IMC-PT. Furthermore, YFCC100M has been often exploited to train deep image matching methods so it can be assumed that some of the compared deep filters are advantaged on this dataset, but this information cannot be retrieved or elaborated. Nevertheless, the proposed modules are handcrafted and not positively biased so the comparison is still valid for the main aim of the evaluation.
The planar dataset contains 135 image pairs from 35 scenes collected from HPatches [hpatches], EVD [mods] and further datasets aggregated from [hess_lapl_affine, dtm, slime]. Each scene usually includes five image pairs except those belonging to EVD consisting of a single image pair. The planar dataset includes scenes with challenging viewpoint changes possibly paired with strong illumination variations. All image pairs are adjusted to be roughly non-upright. Outdoor model weights are preferred for deep modules in the case of planar scenes. A thumbnail gallery for the scenes of the Planar dataset is shown in Fig. 11.
D.3 Error metrics
As discussed in Sec. 4.1.4, it is possible to define a fully metric AUC pose error for IMC-PT dataset and to derive an analogous of the pose error for the planar scenes.
The translation thresholds are reasonable since the 3D scene extensions are roughly of the order of 100 m. This metric pose accuracy estimation provides a reasonable choice able to discriminate when the real baseline is negligible with respect to the whole scene so that the viewpoint change appears as rotation only. In this case, the translation angular component can be quite high as noise but it does not affect the pose estimation effectively. Using Eq. 44 would stress this error without a real justification [heb]. Likewise above, the corresponding AUC values for the chosen pose estimation between Eq. 39 and Eq. 41 will be denoted by AUC, AUC respectively, for the given threshold pair . The final metric accuracy scores are given by the average over the value pairs and indicated AUC and AUC, respectively.
where the average reprojection error in a single image is estimated as
The AUC of the maximum average reprojection error in the common area for each image pair in the dataset and a fixed reprojection threshold
is denoted by AUC. The final homography accuracy is provided as the mean AUC over the different thresholds and is indicated by AUC. The accuracy estimation considering the common area is more reliable than an accuracy estimation by averaging the reprojection error only in the four image corners as in [loftr]. This can be understood by observing that the GT planar homography is a pseudo GT estimated from correspondences, which lie inside the boundaries of both images. As an image point gets reprojected in the other image outside and far from the boundaries, which generally happens with the image corners as depicted in Fig. 2, there is no way to check the real error, which also increases going away from the keypoint matches being employed to estimate the GT homography.
Appendix E More results
E.1 Correlation between error metrics
Figure 13 reports the correlation within the employed error measures for each dataset, assessing the analysis of Sec. 4.2.1. On one hand, after the MAGSAC geometric constraints, for both planar and non-planar scenes there is a high correlation within recall, pose estimation accuracy as AUC, regardless of the estimation is derived from the fundamental or essential matrix, respectively indicated by the ‘’ and ‘’ superscripts, and these both are highly anti-correlated with the percentage of filtered matches. On the other hand, without MAGSAC in the non-planar sceness the recall and the pose accuracy estimated by the essential matrix are highly correlated and both strongly anti-correlated with the number of filtered matches, while precision has an almost moderate correlation with the pose accuracy estimated from the fundamental matrix only. Metric and angular AUC for the same kind of pose estimation have also high correlation levels. For the planar case, without MAGSAC the pose accuracy is only moderately correlated with both precision and recall, as well as anti-correlated with the percentages of the filtered matches within a similar degree.
E.2 Number of matches for each method and dataset
For the sake of completeness, Table 7 reports the average number of input matches for each base method and dataset. This information can be used together with the percentage of filtered matches to get an absolute estimation of the filtered matches. According to the reported values, SuperPoint and ALIKED provide the lowest number of matches followed by SIFT and Key.Net, while DISK, LoFTR and DeDoDe v2 the highest number on matches.
E.3 Further sparse pipelines
E.4 Dense pipelines
E.5 Running times
Overall, according to the running time data, it turns out that the default value =2000 used for the main comparative evaluation in Tables 1-9 is optimal in terms of matching quality only for SIFT and Key.Net, while it slightly decreases the matching results for the other pipelines with respect to =500, which also improves the running times. In detail, moving from =500 to =2000 causes an increment of around 2% of the recall and of about 1% for the AUC pose estimation in the case of SIFT and Key.Net, which contain as initial configuration a higher number of outliers than other pipelines, for which both recall and pose AUC remain instead stable or even worsen. This can be motivated by observing that the next homography discovered in MOP-based approaches is more noisy or even not correct in the long run than the previous ones, and high values force to retrieve a plane and delay the iteration halting. According to these considerations, it is reasonable that forcing the homography retrieval by increasing starts to degrade the found solutions as the initial outlier ratio decreases.
Notice also that MiHo can take part in the MOP early stopping since it introduces further constraints in the homography to be searched, justifying higher running times for MOP than those of MOP+MiHo in some configurations, although MiHo should ideally need more computational resources as two tied homographies are estimated at each step instead of only one. In particular, MOP+MiHo running time equals that of MOP for LoFTR and DeDoDe v2, is greater for SuperPoint, ALIKED, and DISK, and lower for SIFT and Key.Net. Excluding SIFT and Key.Net, MOP+MiHo finds more homographies than MOP, suggesting the reliability of the organization of input matches within the scene as almost-similarity transformations.